# Heavy quark suppression and anisotropic flow at intermediate momentum

### v1

@@ II Heavy quark diffusion and radiation
$\frac{d\Gamma(E_{\bm{p}},k)}{dk}=\frac{g^{2}C_{F}\kappa_{T}(p)}{\pi k}[1+n_{B}(k)][1-n_{F}(E_{\bm{p}-\bm{k}})]\\ \times[(1-x)^{2}+1]\int\frac{d^{2}\bm{p}_{T}}{(2\pi)^{2}}\frac{1}{(\bm{p}_{T}^{2}+m^{2}x^{2}+m_{g}^{2})^{2}}\,,$ (3)
where $\Gamma(E_{\bm{p}},k)$ is the rate for a heavy quark with momentum $\bm{p}$ to emit a gluon with energy $k$. The rate has been found to be proportional to the diffusion coefficient, assuming that gluon has larger transverse momentum than soft momentum transfer. If $m^{2}x^{2}$ is greater than $\bm{p}_{T}^{2}$, gluon emission is suppressed at smaller angle than $\sim m/E_{\bm{p}}$ Dokshitzer:2001zm . In Eq. (1), the radiative energy loss is taken into account as follows:
@@
The heavy quark diffusion coefficient depending on momentum and temperature is important to determine the equilibration rate of heavy quarks in high-temperature QCD plasmas. The temperature dependence comes from running of the coupling constant: $\kappa_{L,T}\propto\alpha_{s}(E_{\bm{p}}T)\alpha_{s}(m_{D}^{2})$ Peigne:2008nd . In gluon-bremsstrahlung, the running coupling constant is determined at the scale, $Q^{2}=(\bm{k}_{T}^{2}+m^{2}x^{2}+m_{g}^{2})/x$ Djordjevic:2013xoa . As the temperature decreases, the coupling becomes stronger and nonperturbative effects start to enter. The momentum diffusion coefficient of a heavy quark grows as the momentum increases. By employing the leading-log momentum dependence Moore:2004tg , the static diffusion coefficient, $\kappa_{L,T}(p=0)$ which involves the nonperturbative effects, becomes the only transport parameter in Eq. (1). The momentum diffusion coefficient has been computed at next-to-leading order Caron-Huot:2007rwy . Since the convergence of the perturbative expansion is poor for a realistic value of the strong coupling, it is desirable to determine it using nonperturbative approaches such as lattice QCD computation Banerjee:2011ra ; Francis:2015daa ; Altenkort:2023eav and phenomenological estimate ALICE:2021rxa ; Xu:2017obm . In this work, the momentum diffusion coefficient is constrained by lattice QCD data for the spatial diffusion coefficient, $D_{s}=2T^{2}/\kappa_{L,T}(p=0)$. Through the lattice QCD data and the running coupling constant, nonperturbative effects can be absorbed at low momenta and temperatures near $T_{c}$.
@@ III Nuclear modification factor and elliptic flow
The anisotropic flow of heavy quarks is generated by asymmetric energy loss in noncentral heavy-ion collisions. A relativistic hydrodynamic simulation provides the time evolution of the spatial distributions of temperature and flow velocity fields which are anisotropic in the transverse plane. Initial conditions, especially the initial geometry and fluctuations of heavy-ion collisions, are a major source of significant uncertainties to determine QGP properties. Since this paper aims to qualitatively study how different mechanisms of energy loss influence the observables, a smooth initial distribution given by an optical Glauber model Miller:2007ri is used. In a background medium generated by hydrodynamic simulations, Eq. (1) is solved with an initial condition where the momentum distribution is given by the differential cross section of heavy meson measured in pp collisions CMS:2017uoy ; ALICE:2019nxm and the spatial distribution is given by the same Glauber model,
@@ III.1 Expanding QGP
In noncentral collisions with nonzero impact parameter, the elliptic deformation of thermal media can be quantified with the spatial eccentricity,
@@
Figure 1 shows typical evolution of the QGP temperature at the center($x=0$, $y=0$) and the spatial eccentricity in noncentral PbPb collisions. They are simulated with (3+1)-dimensional hydrodynamics code MUSIC Schenke:2010nt ; Schenke:2010rr ; Paquet:2015lta using the following hydrodynamic parameters [^1] These parameters have been chosen to produce the maximum temperature similar to that in Ref. Alberico:2013bza . Larger $e_{0}$ should generally be used to create higher-temperature plasmas at greater collision energies.: the initial time, $t_{0}=0.6$ fm, the maximum energy density, $e_{0}=100$ GeV/fm3, the ratio of shear viscosity to entropy density, $\eta/s=0.08$, and the equation of state from lattice QCD data Huovinen:2009yb . In more central collisions with smaller impact parameters, the maximum temperature is higher and the initial spatial eccentricity is lower. While the temperature and the spatial eccentricity decrease with time, the momentum anisotropy is expected to increase. Since heavy quarks are not fully thermalized within the lifetime of QGP, they do not follow the bulk flow of a background medium. In response to an initial spatial deformation, heavy quarks acquire the momentum anisotropy through coupling to the collective motion of the medium. In the following subsections, the profiles in the transverse plane, with the temperature at the center and the spatial eccentricity shown in Fig. 1, are used to calculate the heavy quark flow and the suppression factor at midrapidity.
@@ III.2 Evolution of $R_{AA}$ and $v_{2}$
The nuclear modification factor of heavy quarks indicates the energy-loss effects in QGP, while the elliptic flow reflects how heavy quarks are distributed in the transverse plane and how they are correlated with the flowing thermal media. The anisotropic flow arises due to the asymmetric energy loss: heavy quarks out-of-plane cross longer paths of hot media so they experience more energy loss than those in-plane. Since $R_{AA}^{\rm out}$ is smaller than $R_{AA}^{\rm in}$, heavy quarks obtain positive elliptic flow.
Figure 2 shows the evolution of the suppression factor and the anisotropic flow of charm quarks (with the charm-quark mass $m=1.5$ GeV) for $b=10$ fm. The $R_{AA}$ factor decreases with time and saturate at $t\approx 8$ fm when the plasma reaches the critical point. Meanwhile, $v_{2}$ continuously increases. The elliptic flow is affected by the different path-length as well as the anisotropic temperature and collective flow. In contrast to $R_{AA}$ which decreases mostly in the high-temperature plasmas at early times, $v_{2}$ buildup dominates at late times when the bulk flow is large and the heavy quark coupling with the medium is strong. Similar behavior in the time evolution has been observed previously Rapp:2008qc . Although the interaction rate is high and the flow transfer is efficient in the early stages, it takes time for the flow to accumulate. Due to the development at later times, a significant contribution to $v_{2}$ is also expected at the hadronic phase. Furthermore, $v_{2}$ may be more sensitive than $R_{AA}$ to variations of $T_{c}$ and the uncertainties in the freezeout prescription.
@@ III.4 Bulk flow effect
Given $(t,\bm{x}_{T})$, a hydrodynamic simulation provides the temperature and the flow velocity field which are anisotropic in the transverse plane. For an expanding medium, the local rest frame of the medium is different from the lab frame where $\Delta\bm{x}_{T}=\frac{\bm{p}_{T}}{E_{p}}\Delta t$ is satisfied. $\Delta t^{\prime}$ and $\Delta\bm{x}_{T}^{\prime}$ in the local rest frame are given by $\Delta t^{\prime}=\frac{E_{p}^{\prime}}{E_{p}}\Delta t$ and $\Delta\bm{x}_{T}^{\prime}=\frac{\bm{p}_{T}^{\prime}}{E_{p}}\Delta t$ (where the energies and momenta in two frames are related through the Lorentz transformation with the flow velocity), respectively Baier:2006pt ; Cao:2016gvr ; Ke:2019jbh . Using the flow velocity, the heavy quark distribution is updated according to Eq. (1) in the local rest frame of the medium. The heavy quark diffusion coefficient which determines the energy loss depends on both the temperature and flow velocity of an expanding medium. Thus, the transverse momentum spectrum in an expanding medium is influenced by bulk flow as well as by the asymmetric energy loss along different paths in an elliptically deformed medium. To investigate the contribution of bulk flow to the heavy flavor observables, the $R_{AA}$ and $v_{2}$ have been computed in the anisotropic temperature distribution of plasmas with and without the radial flow.
@@ III.5 Impact parameter and mass effect
For noncentral heavy-ion collisions in Fig. 1, this subsection discusses the heavy flavor observables, focusing on the dependence on impact parameter and mass. The dependence on impact parameter is roughly related to the centrality dependence Broniowski:2001ei . In realistic collisions with event-by-event fluctuations, heavy-ion collisions with the same impact parameter can produce different initial eccentricities. However, this work uses a smooth initial condition given by an optical Glauber model, so the eccentricity and the impact parameter are monotonically correlated.
@@
Figure 5 shows the suppression factor and the elliptic flow of charm quarks with $(2\pi T_{c})D_{s}(T_{c})=2-3$ which are closely aligned with lattice QCD data Banerjee:2011ra ; Francis:2015daa ; Altenkort:2023eav . The heavy quark transport parameter determines the degree of heavy quark energy loss: as $D_{s}$ decreases, both the collisional and radiative energy loss increase. Thus, the nuclear modification factor becomes smaller, and the anisotropic flow increases for smaller $D_{s}$. For fixed $D_{s}$, the $R_{AA}$ factor at $b=3.5$ fm is smaller than at $b=10$ fm because the energy loss in more central collisions is larger due to bigger size and higher temperature of plasmas. On the other hand, the elliptic flow at $b=3.5$ fm is smaller than at $b=10$ fm because the initial eccentricity is greater in more peripheral collisions.
@@
The heavy quark interaction and the coupling to QGP are weaker in more peripheral collisions Nahrgang:2014vza . As the impact parameter increases, the efficiency of transferring an initial geometric deformation to the momentum-space anisotropy decreases. Fig. 8 (b) shows the ratio of the $p_{T}$-integrated $v_{2}$ to the initial spatial eccentricity. The $v_{2}^{i}/\epsilon_{2}$ decreases as the impact parameter increases, and the ratio of bottom quarks is smaller than that of charm quarks.

### v2

@@ II Heavy quark diffusion and radiation
$\frac{d\Gamma(E_{\bm{p}},k)}{dk}=\frac{g^{2}C_{F}\kappa_{T}(p)}{\pi k}[1+n_{B}(K)][1-n_{F}(P-K)]\\ \times[(1-x)^{2}+1]\int\frac{d^{2}\bm{p}_{T}}{(2\pi)^{2}}\frac{1}{(\bm{p}_{T}^{2}+m^{2}x^{2}+m_{g}^{2})^{2}}\,,$ (3)
where $\Gamma(E_{\bm{p}},k)$ is the rate for a heavy quark with momentum $\bm{p}$ to emit a gluon with energy $k$, and $n_{B,F}(K)=[\exp(K\cdot U/T)\mp 1]^{-1}$ (with four-momentum $K=(E_{\bm{k}},\bm{k})$ and fluid velocity $U=(\gamma,\gamma\bm{u})$, where $\gamma=1/\sqrt{1-\bm{u}^{2}}$ is the Lorentz factor) which corresponds to the Bose-Einstein or Fermi-Dirac thermal distribution in the plasma rest frame (for heavy quark, $n_{F}(P-K)$ is negligible). The emission rate has been found to be proportional to the diffusion coefficient, assuming that gluon has larger transverse momentum than soft momentum transfer. If $m^{2}x^{2}$ is greater than $\bm{p}_{T}^{2}$, gluon emission is suppressed at smaller angle than $\sim m/E_{\bm{p}}$ Dokshitzer:2001zm . In Eq. (1), the radiative energy loss is taken into account as follows:
@@
The heavy quark diffusion coefficient depending on momentum and temperature is important to determine the equilibration rate of heavy quarks in high-temperature QCD plasmas. The temperature dependence comes from running of the coupling constant: $\kappa_{L,T}\propto\alpha_{s}(E_{\bm{p}}T)\alpha_{s}(m_{D}^{2})T^{3}$ Peigne:2008nd , where $m_{D}$ is determined by $\ln\left(\frac{m_{D}^{2}}{\Lambda_{\rm QCD}^{2}}\right)=\frac{N_{c}(1+N_{f}/6)}{11N_{c}-2N_{f}}\left(\frac{4\pi T}{m_{D}}\right)^{2}$ Peshier:2006ah . In gluon-bremsstrahlung, the running coupling constant is determined at the scale, $Q^{2}=(\bm{k}_{T}^{2}+m^{2}x^{2}+m_{g}^{2})/x$ Djordjevic:2013xoa . As the temperature decreases, the coupling becomes stronger and nonperturbative effects start to enter. The momentum diffusion coefficient of a heavy quark grows as the momentum increases. By employing the leading-log momentum dependence, $\frac{\kappa_{L}(p)}{\kappa_{L}(p=0)}=\frac{3}{2}\left[\frac{E_{\bm{p}}^{2}}{p^{2}}-\frac{E_{\bm{p}}(E_{\bm{p}}^{2}-p^{2})}{2p^{3}}\ln\frac{E_{\bm{p}}+p}{E_{\bm{p}}-p}\right]$ and $\frac{\kappa_{T}(p)}{\kappa_{T}(p=0)}=\frac{3}{2}\left[\frac{3}{2}-\frac{E_{\bm{p}}^{2}}{2p^{2}}+\frac{(E_{\bm{p}}^{2}-p^{2})^{2}}{4E_{\bm{p}}p^{3}}\ln\frac{E_{\bm{p}}+p}{E_{\bm{p}}-p}\right]$ Moore:2004tg , the static diffusion coefficient, $\kappa_{L,T}(p=0)$ which involves the nonperturbative effects, becomes the only transport parameter in Eq. (1). The momentum diffusion coefficient has been computed at next-to-leading order Caron-Huot:2007rwy . Since the convergence of the perturbative expansion is poor for a realistic value of the strong coupling, it is desirable to determine it using nonperturbative approaches such as lattice QCD computation Banerjee:2011ra ; Francis:2015daa ; Brambilla:2020siz ; Altenkort:2023eav and phenomenological estimate ALICE:2021rxa ; Xu:2017obm . In this work, the momentum diffusion coefficient at $p=0$ and $T=T_{c}$ is constrained by lattice QCD data, while the momentum and temperature dependence are respectively determined by the leading-log result and running coupling. Through the lattice QCD data and the running coupling constant, nonperturbative effects can be absorbed at low momenta and temperatures near $T_{c}$.
@@ III Nuclear modification factor and elliptic flow
The anisotropic flow of heavy quarks is generated by asymmetric energy loss in noncentral heavy-ion collisions. A relativistic hydrodynamic simulation provides the time evolution of the spatial distributions of temperature and flow velocity fields which are anisotropic in the transverse plane. As the temperature and flow velocity enter Eq. (1) through the heavy quark transport coefficients and the gluon emission rate, the Boltzmann equation becomes coupled with a hydrodynamic background. Initial conditions, especially the initial geometry and fluctuations of heavy-ion collisions, are a major source of significant uncertainties to determine QGP properties. Since this paper aims to qualitatively study how different mechanisms of energy loss influence the observables, a smooth initial distribution given by an optical Glauber model Miller:2007ri is used. In a background medium generated by hydrodynamic simulations, Eq. (1) is solved with an initial condition where the momentum distribution is given by the differential cross section of heavy meson measured in pp collisions CMS:2017uoy ; ALICE:2019nxm and the spatial distribution is given by the same Glauber model,
@@ III.1 Expanding QGP
In noncentral collisions with nonzero impact parameter, the elliptic deformation of thermal media can be quantified with the spatial eccentricity Kolb:2001qz ,
@@
Figure 1 shows typical evolution of the QGP temperature at the center($x=0$, $y=0$) and the spatial eccentricity in noncentral PbPb collisions. They are simulated with (3+1)-dimensional hydrodynamics code MUSIC Schenke:2010nt ; Schenke:2010rr ; Paquet:2015lta using the following hydrodynamic parameters [^1] These parameters have been chosen to produce the maximum temperature similar to that in Ref. Alberico:2013bza . Larger $e_{0}$ should generally be used to create higher-temperature plasmas at greater collision energies.: the initial time, $t_{0}=0.6$ fm, the maximum energy density, $e_{0}=100$ GeV/fm3, the ratio of shear viscosity to entropy density, $\eta/s=0.08$, and the equation of state from lattice QCD data Huovinen:2009yb . In more central collisions with smaller impact parameters, the maximum temperature is higher and the initial spatial eccentricity is lower. While the temperature and the spatial eccentricity decrease with time, the momentum anisotropy is expected to increase. Since heavy quarks are not fully thermalized within the lifetime of QGP, they do not follow the bulk flow of a background medium. In response to an initial spatial deformation, heavy quarks acquire the momentum anisotropy by coupling to the collective motion of the medium and asymmetric energy loss. In the following subsections, the temperature and flow velocity fields, which depend on both space and time, are used to calculate the heavy quark flow and the suppression factor at midrapidity.
@@ III.2 Evolution of $R_{AA}$ and $v_{2}$
The nuclear modification factor of heavy quarks indicates the energy-loss effects in QGP, while the elliptic flow reflects how heavy quarks are distributed in the transverse plane and how they are correlated with the flowing thermal media. The anisotropic flow of heavy quark arises from the asymmetric energy loss, coupled with the collective motion of QGP: heavy quarks out-of-plane cross longer paths of hot media so they experience more energy loss than those in-plane. Since $R_{AA}^{\rm out}$ is smaller than $R_{AA}^{\rm in}$, heavy quarks obtain positive elliptic flow.
Figure 2 shows the evolution of the suppression factor and the anisotropic flow of charm quarks in noncentral collisions with $b=10$ fm. The charm-quark mass, $m=1.5$ GeV, and the spatial diffusion coefficient, $(2\pi T_{c})D_{s}(T_{c})=2$ where $D_{s}(T)=2T^{2}/\kappa_{L,T}(p=0,T)$, have been used. The $R_{AA}$ factor decreases with time and saturate at $t\approx 8$ fm when the plasma reaches the hadronization temperature. Meanwhile, $v_{2}$ continuously increases. The elliptic flow is affected by the different path-length as well as the anisotropic temperature and collective flow. In contrast to $R_{AA}$ which decreases mostly in the high-temperature plasmas at early times, $v_{2}$ buildup dominates at late times when the bulk flow is large and the heavy quark coupling with the medium is strong. Similar behavior in the time evolution has been observed previously Rapp:2008qc . Although the interaction rate is high and the flow transfer is efficient in the early stages, it takes time for the flow to accumulate. Due to the development at later times, a significant contribution to $v_{2}$ is also expected at the hadronic phase. Furthermore, $v_{2}$ may be more sensitive than $R_{AA}$ to variations of $T_{c}$ and the uncertainties in the freezeout prescription.
@@ III.3 Collisional and radiative energy loss with $D_{s}$
It should be mentioned that the valid range of the intermediate momentum, where gluon emission from a single scattering is applicable, is unclear. The momentum must be greater than the heavy quark mass but not so high that the LPM effect needs to be considered. Although gluon emission is more involved than photon emission, the LPM effect on the photon emission rate is less than $\sim 30\%$ Arnold:2001ms . If this effect is included in the radiation term, the influence of the total energy loss decreases by $\sim 10\%$. However, the radiative energy loss still affects the $R_{AA}$ and $v_{2}$ in the way discussed above.
@@ III.4 Bulk flow effect
Given $(t,\bm{x}_{T})$, a hydrodynamic simulation provides the temperature and the flow velocity fields which are anisotropic in the transverse plane. For an expanding medium, the local rest frame (with primed coordinates) of the medium is different from the lab frame (with unprimed coordinates) where $\Delta\bm{x}_{T}=\frac{\bm{p}_{T}}{E_{\bm{p}}}\Delta t$ is satisfied. $\Delta t^{\prime}$ and $\Delta\bm{x}_{T}^{\prime}$ in the local rest frame are respectively given by $\Delta t^{\prime}=\frac{E_{\bm{p}}^{\prime}}{E_{\bm{p}}}\Delta t$ and $\Delta\bm{x}_{T}^{\prime}=\frac{\bm{p}_{T}^{\prime}}{E_{\bm{p}}}\Delta t$, where the energies and momenta in two frames are related through the Lorentz transformation: $E_{\bm{p}}^{\prime}=\gamma(E_{\bm{p}}-\bm{u}_{T}\cdot\bm{p}_{T})$ and $\bm{p}_{T}^{\prime}=\bm{p}_{T}+\big{(}-\gamma E_{\bm{p}}+\frac{\gamma^{2}}{1+\gamma}\bm{u}_{T}\cdot\bm{p}_{T}\big{)}\bm{u}_{T}$ Baier:2006pt ; Cao:2016gvr ; Ke:2019jbh . Transforming the heavy quark momentum from the lab frame to the local rest frame and using $\Delta t^{\prime}$ and $\Delta\bm{x}_{T}^{\prime}$ as given above, the heavy quark distribution is updated according to the Boltzmann equation in the local rest frame of the medium. The heavy quark diffusion coefficient which determines the energy loss depends on both the temperature and flow velocity of an expanding medium. Thus, the transverse momentum spectrum in an expanding medium is influenced by bulk flow as well as by the asymmetric energy loss along different paths in an elliptically deformed medium. To investigate the contribution of bulk flow to the heavy flavor observables, the $R_{AA}$ and $v_{2}$ have been computed in the anisotropic temperature distribution of plasmas with and without the radial flow.
@@ III.5 Impact parameter and mass effect
For the noncentral heavy-ion collisions in Sec. III.1, this subsection discusses the heavy flavor observables, focusing on the dependence on impact parameter and mass. The dependence on impact parameter is roughly related to the centrality dependence Broniowski:2001ei . In realistic collisions with event-by-event fluctuations, heavy-ion collisions with the same impact parameter can produce different initial eccentricities. However, this work uses a smooth initial condition given by an optical Glauber model, so the eccentricity and the impact parameter are monotonically correlated.
@@
Figure 5 shows the suppression factor and the elliptic flow of charm quarks with $(2\pi T_{c})D_{s}(T_{c})=2-3$ which are closely aligned with lattice QCD data Banerjee:2011ra ; Francis:2015daa ; Brambilla:2020siz ; Altenkort:2023eav . The heavy quark transport parameter determines the degree of heavy quark energy loss: as $D_{s}$ decreases, both the collisional and radiative energy loss increase. Thus, the nuclear modification factor becomes smaller, and the anisotropic flow increases for smaller $D_{s}$. For fixed $D_{s}$, the $R_{AA}$ factor at $b=3.5$ fm is smaller than at $b=10$ fm because the energy loss in more central collisions is larger due to bigger size and higher temperature of plasmas. On the other hand, the elliptic flow at $b=3.5$ fm is smaller than at $b=10$ fm because the initial eccentricity is greater in more peripheral collisions.
@@
The heavy quark interaction and the coupling to QGP are weaker in more peripheral collisions Nahrgang:2014vza . As the impact parameter increases, the efficiency of transferring an initial geometric deformation to the momentum-space anisotropy decreases. Fig. 8 (b) shows the ratio of the $p_{T}$-integrated $v_{2}$ to the initial spatial eccentricity. The $v_{2}^{\rm int}/\epsilon_{2}$ decreases as the impact parameter increases, and the ratio of bottom quarks is smaller than that of charm quarks.